OpenStudy (anonymous):

the relatioship between the volume of a cube and the lenght of its side is a function. the volume is a fuction of the side lenght. 1)is this a linear function? justify. 2)describe the way in which the gradient of this function changes? 3)what is the average gradient of the function if the length is 2?

OpenStudy (nowhereman):

So what exactly don't you know?

OpenStudy (anonymous):

everything in the question

OpenStudy (nowhereman):

If a cube has all its side length equal to a length x, what is its volume?

OpenStudy (anonymous):

x^3

OpenStudy (nowhereman):

That is what they by the volume function depending on the side length. Next calculate the gradient (=derivative) of this function.

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

is this linear function

OpenStudy (nowhereman):

A linear function f(x) must satisfy the conditions \[f(x+y) = f(x)+f(y)\quad f(αx) = αf(x)\] for any real numbers x,y,α.

OpenStudy (anonymous):

so is this a linear function

OpenStudy (nowhereman):

No, but you should check this yourself.